# Topological chiral superconductivity

### v1

@@ III Chiral superconductors with non-trivial topological orders
Table 1: Topological properties of the chiral superconductors. The table lists the number $N_{\text{top}}$ of topological excitations (with no $A_{\mu}$-flux i.e. no vorticity), the charge-$e_{SF}$ condensation (which determine the minimal electromagnetic $A_{\mu}$-flux quantum), chiral central charge $c$ (i.e. the number of chiral edge modes), the fractions for each spin-valley components $f_{I}$, the topological Hall response $\sigma=\boldsymbol{e}^{\text{top}\top}(K^{\text{top}})^{-1}\boldsymbol{e}^{\text{top}}$, the average orbital angular momentum $\langle L\rangle$ per electron, and the type of topological order in the topological chiral superconductor.
$K$-matrix $N_{\text{top}}$ $e_{SF}$ $c$ $f_{I}$ $\sigma$ $\langle L\rangle$ type of topological order
@@
$K_{2b}$ (32) 3 $2$ 1 $\frac{1}{2},\frac{1}{2}$ $\frac{1}{3}$ $-\frac{3}{2}$ $K^{\text{top}}=(3)$, $\boldsymbol{e}^{\text{top}}=(1)$
$K_{2c}$ (33) 5 $2$ 1 $\frac{1}{2},\frac{1}{2}$ $\frac{1}{5}$ $-\frac{5}{2}$ $K^{\text{top}}=(5)$, $\boldsymbol{e}^{\text{top}}=(1)$
$K_{3a}$ (34) 1 $4$ 0 $\frac{1}{4},\frac{1}{2},\frac{1}{4}$ 00 $1$ $K^{\text{top}}=\begin{pmatrix}-1&0\\ 0&1\\ \end{pmatrix}$, $\boldsymbol{e}^{\text{top}}=(1,1)$
$K_{3b}$ (35) 1 $6$ 0 $\frac{1}{6},\frac{1}{2},\frac{1}{3}$ $-8$ $-1$ $K^{\text{top}}=\begin{pmatrix}-1&0\\ 0&1\\ \end{pmatrix}$, $\boldsymbol{e}^{\text{top}}=(-3,1)$
[…]
$K_{4a}$ (37) 1 $4$ $-1$ $\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}$ $-25$ $-\frac{1}{2}$ $K^{\text{top}}=\begin{pmatrix}-1&0&0\\ 0&-1&0\\ 0&0&1\\ \end{pmatrix}$, $\boldsymbol{e}^{\text{top}}=(1,7,5)$
$K_{4b}$ (38) 12 $4$ 3 $\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}$ $-1$ $-\frac{1}{2}$ $K^{\text{top}}=\begin{pmatrix}-2&0&0\\ 0&3&0\\ 0&0&-2\\ \end{pmatrix}$, $\boldsymbol{e}^{\text{top}}=(2,3,2)$
$K_{4c}$ (39) 1 $4$1 $\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}$ $-1$ $\frac{1}{2}$ $K^{\text{top}}=\begin{pmatrix}-1&0&0\\ 0&-1&0\\ 0&0&1\\ \end{pmatrix}$, $\boldsymbol{e}^{\text{top}}=(3,1,3)$
The above chiral superconductors described by (II) usually have non-trivial topological orders and are beyond BCS. We now compute the topological order and other properties of such chiral superconductors. Following Ref. 31, 8, 12, we start with the effective Lagrangian for the chiral superconductor described by (II):
@@
Such a matrix is symmetric and integral. Since $D_{n\kappa}=0$, the last column of $\widetilde{K}$ is zero. The last row of $\widetilde{K}$ is also zero since $\widetilde{K}$ is symmetric.
When there is a gapless mode, we must include Maxwell terms as the new leading order contribution in the superfluid sector. This includes a term $g_{SF}\widetilde{f}_{SF,\mu\nu}\widetilde{f}_{SF}^{\mu\nu}$, where $\widetilde{f}_{SF,\mu\nu}=\partial_{\mu}\widetilde{a}_{SF,\nu}-\partial_{\nu}\widetilde{a}_{SF,\mu}$, and $g_{I}\widetilde{f}_{SF,\mu\nu}\widetilde{f}_{I}^{\mu\nu}$ where $\widetilde{f}_{I,\mu\nu}=\partial_{\mu}\widetilde{a}_{I,\nu}-\partial_{\nu}\widetilde{a}_{I,\mu}$. Writing ^$\hat{\partial}^{\mu\nu}=\epsilon^{\mu\nu\lambda}\partial_{\lambda}$, the effective Lagrangian (III) becomes
${\cal L}=$ $\;\widetilde{\boldsymbol{a}}_{\mu}\begin{pmatrix}\tfrac{1}{4\pi}K^{\text{top}}\;\hat{\partial}^{\mu\nu}&\boldsymbol{g}\;{\hat{\partial}^{\mu}}_{\sigma}\hat{\partial}^{\sigma\nu}\\ \boldsymbol{g}^{\top}\;{\hat{\partial}^{\mu}}_{\sigma}\hat{\partial}^{\sigma\nu}&g_{SF}\;{\hat{\partial}^{\mu}}_{\sigma}\hat{\partial}^{\sigma\nu}&\end{pmatrix}\widetilde{\boldsymbol{a}}_{\nu}+\frac{\widetilde{\boldsymbol{e}}}{2\pi}A_{\mu}\hat{\partial}^{\mu\nu}\widetilde{\boldsymbol{a}}_{\nu}$
@@
The basis of (45) is not unique, and there is an ambiguity in separating out the superfluid mode. In particular, we can add a multiple of $a_{SF}$ onto any of the topological gauge fields and preserve the form of 45. In the new basis, the $g_{I}$ have changed. Had we included Maxwell terms in the topological sector, a full matrix $g_{\text{tot}}$ would transform under this basis change, but we continue to throw out any mixing of $g_{I}$ into the purely topological sector. Any lowest order terms in our analysis will not be affected by this truncation.
@@
${\cal L}^{EM}=$ $\;\frac{1}{4}\hat{\partial}^{\mu\lambda}A_{\lambda}\frac{\widetilde{\boldsymbol{e}}^{\top}}{2\pi}M_{\mu\nu}^{-1}\frac{\widetilde{\boldsymbol{e}}}{2\pi}\hat{\partial}^{\nu\rho}A_{\rho}$ (51)
$M^{\mu\nu}=$ $\;\begin{pmatrix}\tfrac{1}{4\pi}K^{\text{top}}\;\hat{\partial}^{\mu\nu}+i\epsilon&\boldsymbol{g}\;{\hat{\partial}^{\mu}}_{\sigma}\hat{\partial}^{\sigma\nu}\\ \boldsymbol{g}^{\top}\;{\hat{\partial}^{\mu}}_{\sigma}\hat{\partial}^{\sigma\nu}&g_{SF}\;{\hat{\partial}^{\mu}}_{\sigma}\hat{\partial}^{\sigma\nu}+i\epsilon\end{pmatrix}$ (52)
where we added an $i\epsilon$ identity term to the quadratic-in-$\widetilde{\boldsymbol{a}}$ term to make the integrals converge. The inverse of $M$ can be written in terms of $g$ and $K$ and their inverses, and these can be expanded order by order in derivatives.
The lowest order term is
$M_{SF}^{-1}=\begin{pmatrix}0&\boldsymbol{0}^{\top}\\ \boldsymbol{0}&\left(g_{SF}\;{\hat{\partial}^{\mu}}_{\sigma}\hat{\partial}^{\sigma\nu}+i\epsilon\right)^{-1}\end{pmatrix}$ (53)
Some care must be taken combining the two inverse derivatives with the derivatives on $A_{\mu}$. We work in Lorentz signature (,+,+)(-,+,+)( - , + , + ), Fourier transform, and fix the gauge to $a_{0}=0$. In a basis $\{\widetilde{a}_{t},\widetilde{a}_{\ell}\}=\{(0,-p_{y},p_{x}),(0,p_{x},p_{y})\}$ labeling the transverse and longitudinal components respectively, we obtain
@@
Returning to momentum space, a general vector potential $\widetilde{A}$ can be separated into longitudinal and transverse parts via the projectors
$\widetilde{A}_{\ell}^{i}=\frac{p^{i}p_{j}}{p_{x}^{2}+p_{y}^{2}}\widetilde{A}^{j},\quad\widetilde{A}_{t}^{i}=\frac{\epsilon^{i\ell}p_{\ell}\epsilon_{jk}p^{k}}{p_{x}^{2}+p_{y}^{2}}\widetilde{A}^{j}$ (54)
@@
${\cal L}^{EM}_{SF}=\frac{\widetilde{e}_{SF}^{2}}{4(2\pi)^{2}g_{SF}}$ $\left(\widetilde{A}^{\mu}_{t}\widetilde{A}_{t,\mu}+\frac{\omega^{2}+p_{x}^{2}+p_{y}^{2}}{\omega^{2}-p_{x}^{2}-p_{y}^{2}+i\epsilon}\widetilde{A}_{\ell}^{\mu}\widetilde{A}_{\ell,\mu}\right)$ (55)
@@
$J_{SF}^{\mu}$ $\;=\frac{\widetilde{e}_{SF}^{2}}{2(2\pi)^{2}g_{SF}}\left(\frac{\epsilon^{\mu i}p_{i}\epsilon_{jk}p^{k}}{p_{x}^{2}+p_{y}^{2}}\right.$
$+\left.\left(\frac{\omega^{2}+p_{x}^{2}+p_{y}^{2}}{\omega^{2}-p_{x}^{2}-p_{y}^{2}+i\epsilon}\right)\frac{p^{\mu}p_{j}}{p_{x}^{2}+p_{y}^{2}}\right)\widetilde{A}^{j}$ (56)
@@
The next order contribution to $M^{-1}$ describes the Hall response of the topological sector.
$M^{-1}_{top}=4\pi\begin{pmatrix}(K^{\text{top}}\hat{\partial}^{\mu\nu}+i\epsilon)^{-1}&\frac{-(K^{\text{top}}\hat{\partial}^{\mu\nu}+i\epsilon)^{-1}\boldsymbol{g}}{g_{SF}}\\ \frac{-\boldsymbol{g}^{\top}(K^{\text{top}}\hat{\partial}^{\mu\nu}+i\epsilon)^{-1}}{g_{SF}}&\frac{\boldsymbol{g}^{\top}(K^{\text{top}}\hat{\partial}^{\mu\nu}+i\epsilon)^{-1}\boldsymbol{g}}{g_{SF}^{2}}\end{pmatrix}$ (57)
Contracting with ^$\hat{\partial}^{\mu\nu}A_{\nu}$, factors of ^$\hat{\partial}^{\mu\nu}$ cancel without poles, so we take $\epsilon$ to zero. We obtain the Chern-Simons factor in the EM Lagrangian
${\cal L}^{EM}_{top}=\frac{\widetilde{\boldsymbol{e}}^{\top}}{4\pi}\begin{pmatrix}(K^{\text{top}})^{-1}&\frac{-(K^{\text{top}})^{-1}\boldsymbol{g}}{g_{SF}}\\ \frac{-\boldsymbol{g}^{\top}(K^{\text{top}})^{-1}}{g_{SF}}&\frac{\boldsymbol{g}^{\top}(K^{\text{top}})^{-1}\boldsymbol{g}}{g_{SF}^{2}}\end{pmatrix}\widetilde{\boldsymbol{e}}\;\epsilon^{\mu\nu\lambda}A_{\mu}\partial_{\nu}A_{\lambda}$ (58)
where $\epsilon^{\prime}$ carries a factor of $\hat{p}^{-1}$. From this the Hall response can be written as
$J_{top}^{\mu}=\frac{\widetilde{\boldsymbol{e}}^{\top}}{2\pi}\begin{pmatrix}(K^{\text{top}})^{-1}&\frac{-(K^{\text{top}})^{-1}\boldsymbol{g}}{g_{SF}}\\ \frac{-\boldsymbol{g}^{\top}(K^{\text{top}})^{-1}}{g_{SF}}&\frac{\boldsymbol{g}^{\top}(K^{\text{top}})^{-1}\boldsymbol{g}}{g_{SF}^{2}}\end{pmatrix}\widetilde{\boldsymbol{e}}\;\epsilon^{\mu\nu\lambda}\partial_{\nu}A_{\lambda}$ (59)
This term includes $\boldsymbol{g}$, but one can check it is invariant under a basis change redefining the topological sector relative to the superfluid component. In fact, there is such a basis change in which $\boldsymbol{g}=\boldsymbol{0}$, in which case the Hall response reduces to
$J_{top}^{\mu}$ $=\sigma\epsilon^{\mu\nu\lambda}\partial_{\nu}A_{\lambda}$
$\sigma$ $\equiv e_{I}^{\text{top}}(K^{\text{top}})^{-1}_{IJ}e_{J}^{\text{top}}$ (60)
In this manner, the matrix $g_{\text{tot}}$ acts as a metric on the gauge fields. The basis in which the topological response separates from the superfluid response is the basis in which $\widetilde{a}_{I}g_{\text{tot}}\widetilde{a}_{SF}=0$. This basis may not preserve the vortex quantization, but it does preserve quantization of the Hall response. We can see this explicitly by noting that from a vortex-quantized basis, $\boldsymbol{g}$ becomes zero under the transformation $\widetilde{a}_{I}\rightarrow\widetilde{a}_{I}+t_{I}\widetilde{a}_{SF}$ where $t_{I}=-g_{I}/g_{SF}$. $\widetilde{\boldsymbol{e}}$ transforms as the dual, $\widetilde{e}_{I}\rightarrow\widetilde{e}_{I}$ and $\widetilde{e}_{SF}\rightarrow\widetilde{e}_{SF}+t_{I}\widetilde{e}_{I}$. Since this transformation only affects $\widetilde{e}_{SF}$, it does not change the universal Hall response in III. Thus the total response current is
$J^{i}$ $\;=\sigma\epsilon^{ij}\omega\widetilde{A}_{j}+\frac{\widetilde{e}_{SF}^{2}}{2(2\pi)^{2}g_{SF}}\left(\frac{\epsilon^{i\ell}p_{\ell}\epsilon_{jk}p^{k}}{p_{x}^{2}+p_{y}^{2}}\right.$
$+\left.\left(\frac{\omega^{2}+p_{x}^{2}+p_{y}^{2}}{\omega^{2}-p_{x}^{2}-p_{y}^{2}+i\epsilon}\right)\frac{p^{i}p_{j}}{p_{x}^{2}+p_{y}^{2}}\right)\widetilde{A}^{j}$ (61)
where $\omega\widetilde{A}_{j}$ is the electric field, and as before we use Latin indices to indicate spatial degrees of freedom. We see that the topological electromagnetic response $\sigma\epsilon^{ij}\omega\widetilde{A}_{j}$ can be separated from other electromagnetic responses, and we can measure the topological Hall conductance $\sigma$ even in the superfluid phase.
To obtain the topological order in the chiral superconductor, we need to study its topological excitations, which are labeled by integer vectors $(\widetilde{l}_{I},\widetilde{l}_{SF})$. The term $-\widetilde{a}_{I0}\widetilde{l}_{I}\delta(\boldsymbol{x})-\widetilde{a}_{SF,0}\widetilde{l}_{SF}\delta(\boldsymbol{x})$ in the effective Lagrangian describes such a topological excitation located at $\boldsymbol{x}=0$. From the equation of motion for $\widetilde{a}_{SF,0}$,
@@
We see that if $\widetilde{l}_{SF}\neq 0$, the excitation corresponds to a vortex in the chiral superconductor [8, 12]. Such a vortex carries a $\widetilde{e}_{SF}^{-1}$ unit of $A_{\mu}$-flux:
@@
When $\widetilde{l}_{SF}=0$, the excitations labeled by $\widetilde{\boldsymbol{l}}^{\top}=(\widetilde{l}_{I})$ are finite energy excitations in the chiral superconductor (i.e. not vortices). This kind of excitations has statistics and mutual statistics
$s_{\widetilde{\boldsymbol{l}}}=\frac{1}{2}\widetilde{\boldsymbol{l}}^{\top}(K^{\text{top}})^{-1}\widetilde{\boldsymbol{l}},\ \ \ \ s_{\widetilde{\boldsymbol{l}},\widetilde{\boldsymbol{l}}^{\prime}}=\widetilde{\boldsymbol{l}}^{\top}(K^{\text{top}})^{-1}\widetilde{\boldsymbol{l}}^{\prime}.$ (65)
@@
$e_{SF}$ $=2,\ \ \sigma=\frac{1}{3},\ \ {\boldsymbol{e}^{\text{top}}}=(1)$ (69)

### v3

@@ III Chiral superconductors with non-trivial topological orders
Table 1: Topological properties of the chiral superconductors. The table lists the number $N_{\text{top}}$ of topological excitations (with no $A_{\mu}$-flux i.e. no vortitopologicalcity), the charge-$e_{SF}$ condensation (which determine the minimal electromagnetic $A_{\mu}$-flux quantum), chiral central charge $c$ (i.e. the number of chiral edge modes), the fractions for each spin-valley components $f_{I}$, the average orbital angular momentum $\langle L\rangle$ per electron, and the type of topological order in the topological chiral superconductor.
$K$-matrix $N_{\text{top}}$ $e_{SF}$ $c$ $f_{I}$ $\langle L\rangle$ type of topological order
@@
$K_{2b}$ (32) 3 $2$ 1 $\frac{1}{2},\frac{1}{2}$ $-\frac{3}{2}$ $K^{\text{top}}=(3)$, $\boldsymbol{e}^{\text{top}}=(1)$
$K_{2c}$ (33) 5 $2$ 1 $\frac{1}{2},\frac{1}{2}$ $-\frac{5}{2}$ $K^{\text{top}}=(5)$, $\boldsymbol{e}^{\text{top}}=(1)$
$K_{3a}$ (34) 1 $4$ 0 $\frac{1}{4},\frac{1}{2},\frac{1}{4}$ $1$ $K^{\text{top}}=\begin{pmatrix}-1&0\\ 0&1\\ \end{pmatrix}$, $\boldsymbol{e}^{\text{top}}=(1,1)$
$K_{3b}$ (35) 1 $6$ 0 $\frac{1}{6},\frac{1}{2},\frac{1}{3}$ $-1$ $K^{\text{top}}=\begin{pmatrix}-1&0\\ 0&1\\ \end{pmatrix}$, $\boldsymbol{e}^{\text{top}}=(-3,1)$
$K_{3c}$ (36) 10 $4$ 0 $\frac{1}{4},\frac{1}{2},\frac{1}{4}$ $-1$ $K^{\text{top}}=\begin{pmatrix}-5&0\\ 0&2\\ \end{pmatrix}$, $\boldsymbol{e}^{\text{top}}=(9,4)$
$K_{4a}$ (37) 1 $4$ $-1$ $\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}$ $-\frac{1}{2}$ $K^{\text{top}}=\begin{pmatrix}-1&0&0\\ 0&-1&0\\ 0&0&1\\ \end{pmatrix}$, $\boldsymbol{e}^{\text{top}}=(1,7,5)$
$K_{4b}$ (38) 12 $4$ 3 $\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}$ $-\frac{1}{2}$ $K^{\text{top}}=\begin{pmatrix}-2&0&0\\ 0&3&0\\ 0&0&-2\\ \end{pmatrix}$, $\boldsymbol{e}^{\text{top}}=(2,3,2)$
$K_{4c}$ (39) 1 $4$1 $\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}$ $\frac{1}{2}$ $K^{\text{top}}=\begin{pmatrix}-1&0&0\\ 0&-1&0\\ 0&0&1\\ \end{pmatrix}$, $\boldsymbol{e}^{\text{top}}=(3,1,3)$
The above chiral superconductors described by (II) usually have non-trivial topological orders and are beyond BCS. We now compute the topological order and other properties of such chiral superconductors. Following Ref. 36, 8, 12, we start with the effective Lagrangian for the chiral superconductor described by (II):
@@
Such a matrix is symmetric and integral. Since $D_{I\kappa}=0$, the last column of $\widetilde{K}$ is zero. The last row of $\widetilde{K}$ is also zero since $\widetilde{K}$ is symmetric.
When there is a gapless mode, we must include Maxwell terms as the new leading order contribution in the superfluid sector. Most generally, this includes terms $g_{IJ}\widetilde{f}_{I,\mu\nu}\widetilde{f}_{J}^{\mu\nu}$, where $\widetilde{f}_{I,\mu\nu}=\partial_{\mu}\widetilde{a}_{I,\nu}-\partial_{\nu}\widetilde{a}_{I,\mu}$. We further decompose $g$ into $g^{\text{top}}$ purely in the topological sector, $\boldsymbol{g}$ coupling the topological and superfluid sectors, and $g_{SF}$ purely in the superfluid sector.
Fourier transforming and writing $p^{\mu\nu}=\epsilon^{\mu\nu\lambda}p_{\lambda}$, the effective Lagrangian (III) becomes at leading order
${\cal L}=\frac{\widetilde{\boldsymbol{e}}}{2\pi}A_{\mu}p^{\mu\nu}\widetilde{\boldsymbol{a}}_{\nu}-\widetilde{\boldsymbol{l}}\cdot\widetilde{\boldsymbol{a}}\;\delta(\boldsymbol{x})$ (49)
$+\widetilde{\boldsymbol{a}}_{\mu}\begin{pmatrix}\tfrac{K^{\text{top}}}{4\pi}\;p^{\mu\nu}+g^{\text{top}}\;{p^{\mu}}_{\sigma}p^{\sigma\nu}&\boldsymbol{g}\;{p^{\mu}}_{\sigma}p^{\sigma\nu}\\ \boldsymbol{g}\;{p^{\mu}}_{\sigma}p^{\sigma\nu}&g_{SF}\;{p^{\mu}}_{\sigma}p^{\sigma\nu}&\end{pmatrix}\widetilde{\boldsymbol{a}}_{\nu}$
The splitting of the basis into a superfluid and topological sector is not well-defined, and in particular we are free to add the superfluid mode to any topological basis vector (say with coefficients $t_{I^{\prime}}$, where primed indices go up to $\kappa-1$) and preserve the form of $K^{\text{top}}$. Under such a transformation, the gauge field labels transform as $a_{SF}\rightarrow a_{SF}-t_{I^{\prime}}a_{I^{\prime}}$, the charge vector transforms as $e_{I^{\prime}}\rightarrow e_{I^{\prime}}+t_{I^{\prime}}e_{SF}$, the excitation label transforms as $\widetilde{l}_{I^{\prime}}\rightarrow\widetilde{l}_{I^{\prime}}+t_{I^{\prime}}\widetilde{l}_{SF}$, and the $g_{I^{\prime}}$ transform as $g_{I^{\prime}}+t_{I^{\prime}}g_{SF}$. Therefore we can block-diagonalize the $g_{IJ}$ by choosing $t_{I^{\prime}}=-g_{I^{\prime}}/g_{SF}$. This basis is no longer charge quantized, but the superfluid mode is now decoupled from the topological sector.
${\cal L}=$ $\;\widetilde{\boldsymbol{a}}_{\mu}\begin{pmatrix}\tfrac{1}{4\pi}K^{\text{top}}\;p^{\mu\nu}&\boldsymbol{0}\\ \boldsymbol{0}^{\top}&g_{SF}\;{p^{\mu}}_{\sigma}p^{\sigma\nu}&\end{pmatrix}\widetilde{\boldsymbol{a}}_{\nu}+\frac{\widetilde{\boldsymbol{e}}}{2\pi}A_{\mu}\hat{p}^{\mu\nu}\widetilde{\boldsymbol{a}}_{\nu}$
@@
We have also dropped $g^{\text{top}}$ as a subleading contribution that will not contribute at any order relevant in the transport calculations which follow. We stress that now $\widetilde{\boldsymbol{e}}=(\widetilde{\boldsymbol{e}}^{\text{top}},e_{SF})$ with $\widetilde{\boldsymbol{e}}^{\text{top}}$ no longer quantized as integers, but $e_{SF}$ is unchanged and is still an integer.
@@
${\cal L}^{EM}$ $\;=\frac{1}{4}p^{\mu\lambda}A_{\lambda}\frac{\widetilde{e}^{\text{top}}_{I}}{2\pi}\left(\tfrac{1}{4\pi}K^{\text{top}}p_{\mu\nu}+i\epsilon\right)^{-1}_{IJ}\frac{\widetilde{e}^{\text{top}}_{J}}{2\pi}p^{\nu\rho}A_{\rho}$
$+\frac{e^{2}_{SF}}{4(2\pi)^{2}}p^{\mu\lambda}A_{\lambda}\left(g_{SF}\;{\hat{\partial}^{\mu}}_{\sigma}\hat{\partial}^{\sigma\nu}+i\epsilon\right)^{-1}p^{\nu\rho}A_{\rho}$ (52)
where we added an $i\epsilon$ identity term to the quadratic-in-$\widetilde{\boldsymbol{a}}$ term to make the integrals converge.
In the first term, a $p^{\mu\lambda}$ and $(p_{\mu\nu})^{-1}$ cancel each other without poles, so we can simplify to find the transverse conductance
${\cal L}^{EM}_{\text{top}}=\sigma\epsilon^{\lambda\mu\nu}A_{\lambda}\partial_{\mu}A_{\nu}$
$\sigma=\frac{1}{4\pi}\widetilde{e}^{\text{top}}_{I}\left(K^{\text{top}}\right)^{-1}_{IJ}\widetilde{e}^{\text{top}}_{J}$ (53)
To emphasize, this Hall conductance is not quantized because the $\widetilde{e}^{\text{top}}$ chosen to block diagonalize $g_{IJ}$ are not integer.
Some care must be taken contracting the momenta in the superfluid term. We work in Lorentz signature (,+,+)(-,+,+)( - , + , + ), Fourier transform, and fix the gauge to $a_{0}=0$. In a basis $\{\widetilde{a}_{t},\widetilde{a}_{\ell}\}=\{(0,-p_{y},p_{x}),(0,p_{x},p_{y})\}$ labeling the transverse and longitudinal components respectively, we obtain
@@
Returning to momentum space, a general vector potential $A_{\mu}$ can be separated into longitudinal and transverse parts via the projectors
$A_{\ell}^{i}=\frac{p^{i}p_{j}}{p_{x}^{2}+p_{y}^{2}}A^{j},\quad A_{t}^{i}=\frac{\epsilon^{i\ell}p_{\ell}\epsilon_{jk}p^{k}}{p_{x}^{2}+p_{y}^{2}}A^{j}$ (54)
@@
${\cal L}^{EM}_{SF}=\frac{e_{SF}^{2}}{4(2\pi)^{2}g_{SF}}$ $\left(A^{\mu}_{t}A_{t,\mu}+\frac{\omega^{2}+p_{x}^{2}+p_{y}^{2}}{\omega^{2}-p_{x}^{2}-p_{y}^{2}+i\epsilon}A_{\ell}^{\mu}A_{\ell,\mu}\right)$ (55)
@@
$J_{SF}^{\mu}$ $\;=\frac{e_{SF}^{2}}{2(2\pi)^{2}g_{SF}}\left(\frac{\epsilon^{\mu i}p_{i}\epsilon_{jk}p^{k}}{p_{x}^{2}+p_{y}^{2}}\right.$
$\ \ \ \ \ +\left.\frac{\omega^{2}+p_{x}^{2}+p_{y}^{2}}{\omega^{2}-p_{x}^{2}-p_{y}^{2}+i\epsilon}\ \frac{p^{\mu}p_{j}}{p_{x}^{2}+p_{y}^{2}}\right)A^{j}$ (56)
@@
Thus the total response current is
$J^{i}$ $\;=\sigma\epsilon^{ij}\omega A_{j}+\frac{e_{SF}^{2}}{2(2\pi)^{2}g_{SF}}\left(\frac{\epsilon^{i\ell}p_{\ell}\epsilon_{jk}p^{k}}{p_{x}^{2}+p_{y}^{2}}\right.$
$\ \ \ \ \ +\left.\frac{\omega^{2}+p_{x}^{2}+p_{y}^{2}}{\omega^{2}-p_{x}^{2}-p_{y}^{2}+i\epsilon}\ \frac{p^{i}p_{j}}{p_{x}^{2}+p_{y}^{2}}\right)A^{j}$ (57)
where $\omega A_{j}$ is the electric field, and as before we use Latin indices to indicate spatial degrees of freedom.
To obtain the topological order in the chiral superconductor, let us still use the basis in the effective Lagrangian (III), where the superfluid sector and the topological sector completely decouple. We need to study its topological excitations, which are labeled by vectors $(\widetilde{l}_{I^{\prime}},\widetilde{l}_{SF})$, $I^{\prime}=1,\cdots,\kappa-1$. In this basis, $\widetilde{l}_{SF}$ is still an integer, but $\widetilde{l}_{I^{\prime}}$ may not be integer. However, when $\widetilde{l}_{SF}=0$, $\widetilde{l}_{I^{\prime}}$ are quantized as integers.
The term $-\widetilde{a}_{I0}\widetilde{l}_{I}\delta(\boldsymbol{x})-\widetilde{a}_{SF,0}\widetilde{l}_{SF}\delta(\boldsymbol{x})$ in the effective Lagrangian describes such a topological excitation located at $\boldsymbol{x}=0$. From the equation of motion for $\widetilde{a}_{SF,0}$,
@@
We see that if $\widetilde{l}_{SF}\neq 0$, the excitation corresponds to a vortex in the chiral superconductor [8, 12]. Such a vortex carries a $e_{SF}^{-1}$ unit of $A_{\mu}$-flux:
@@
When $\widetilde{l}_{SF}=0$, the excitations labeled by integers $\widetilde{\boldsymbol{l}}^{\text{top},\top}=(\widetilde{l}_{I^{\prime}})$ are finite energy excitations in the chiral superconductor (i.e. not vortices). This kind of excitations carry fractional electric charges
$q_{\widetilde{\boldsymbol{l}}^{\text{top}}}=\widetilde{\boldsymbol{e}}_{1}^{\text{top}\top}(K^{\text{top}})^{-1}\widetilde{\boldsymbol{l}}^{\text{top}}.$ (61)
which are not quantized. They also have statistics and mutual statistics
$s_{\widetilde{\boldsymbol{l}}^{\text{top}}}$ $=\frac{1}{2}\widetilde{\boldsymbol{l}}^{\text{top}\top}(K^{\text{top}})^{-1}\widetilde{\boldsymbol{l}}^{\text{top}},$
$s_{\widetilde{\boldsymbol{l}}_{1}^{\text{top}},\widetilde{\boldsymbol{l}}_{2}^{\text{top}}}$ $=\widetilde{\boldsymbol{l}}_{1}^{\text{top}\top}(K^{\text{top}})^{-1}\widetilde{\boldsymbol{l}}_{2}^{\text{top}}.$ (62)
@@
$e_{SF}$ $=2,\ \ \ \ {\boldsymbol{e}^{\text{top}}}=(1)$ (66)